Suppose the market for agrochemicals has a supply




Sx=30+A and the demand curve of Dq=240-2q assume that the market is perfectly competitive what would the equilibrium price and quantity be?

1
Expert's answer
2021-11-28T17:49:00-0500

Solution:

At equilibrium: Dq = Sx

240 – 2q = 30 + A

240 – 30 = A + 2q

210 = A + 2q

Divide both sides by 2:

105 = "\frac{A}{2}" + q


q = 105 – "\frac{A}{2}"

Equilibrium quantity = 105 – "\frac{A}{2}"


Substitute in the supply function:

Sx = 30 + A

Sx = 30 + (105 – "\frac{A}{2}")

A = 270 – 2S

Equilibrium price = 270 – 2S